OpenStudy (anonymous):

4-t=3(t-1)-5 A. 3 B. 6 C. No solution D. Indentity

OpenStudy (anonymous):

A

OpenStudy (anonymous):

4-t=3(t-1)-5 4-t=3t-3-5 4-t=3t-8 4=4t-8 12=4t 3=t